An object has a mass of 2 kg is dropped from the top of building 30 meters tall. The initial velocity is zero. As it falls, the object encounters air resistance that is equal to 1/3 v. What is the velocity and altitude of the object after 1.5 seconds?


1
Expert's answer
2021-12-02T00:54:28-0500
"2\frac{dv}{dt}=2(9.8)-\frac{1}{3}v\\58.8(1-e^{-0.166667t})\\
v'=13\frac{m}{s}\\h'=19.8\ m"